OpenStudy (anonymous):

A bread recipe calls for 2/3 cup of white flour and 1 1/2 cups of whole wheat flour. What is the total amount of flour needed

OpenStudy (anonymous):

2/3 + 3/2= 4/6 + 9/6= 13/6

OpenStudy (anonymous):

a. 1 cups b. 1 cups c. 2 cups d. 1 cups . one of these

OpenStudy (anonymous):

1 2/5 cups or 1 3/5 cups or 2 1/6 cups or 1 1/2 cup

OpenStudy (anonymous):

1 3/6

OpenStudy (anonymous):

no it can be its 1 2/5 cups or 1 3/5 cups or 2 1/6 cups or 1 1/2 cup